# Are Vision Language Models Texture or Shape Biased and Can We Steer Them?

### v1

@@ 4.1 Main Results
The results in Fig. 2 (a full table is shown in the appendix) paint a fairly uniform picture. Across different models and on two different tasks, most VLMs seem to perform relatively similarly in terms of shape bias. Shape bias of VLMs is still significantly lower than in humans ($96\%$), but overall higher than in typical image-only discriminative classifiers (e.g., $22\%$ for an ImageNet-trained ResNet-50 [59, 1]). For most models, shape bias is higher than the shape bias of CLIP ViT-L/14 [11] ($60\%$) which is a common vision encoder used in many of our tested models.
On average, we observe a slightly higher shape bias for the image captioning task at some cost in accuracy (mostly due to generic captions that do not mention any objects) and vice-versa. For image description tasks, it can be significantly lower ($38.2\%$) for some outliers but generally ranges from $54.1$ $73.2\%$. For VQA it ranges from $52.9$ $73.8\%$. In this task, we had to remove some models as they did not follow instructions. Some of these models are on the lower end of the shape bias in captioning, which raises the question of whether there might exist a correlation between instruction tuning quality and shape bias (to answer this question, more samples would be needed for a conclusive answer).
Which models are the most shape-biased?
The strongest shape bias is observed in InstructBLIP Vicuna-7B [23] for VQA, but the model generally shows a lower accuracy compared to other models. A more accurate model is InternVL-Chat 1.1 [2] which ranks second place for VQA but first for captioning.
GPT-4V is an outlier.
Given that GPT-4V often achieves SOTA performance and is considered an important baseline, it has a surprisingly poor accuracy compared to most other models - mostly due to refusal to answer: 131/1280 VQA conversations, i.e., roughly 10% which is substantially higher than the refusal rate of all other models ($<1\%$). Note that refusal rates do not affect the shape bias measurement. GPT-4V is also the model with the largest amount of generic image captions ($60.4\%$). We acknowledge that other prompts may have led to other results, however, the result is noteworthy, as the other VLMs mostly behave well under the same prompts. Interestingly, GPT-4V is also significantly more texture-biased than most models in both tasks.
Does scale matter?
LLM capacity does not seem to correlate with shape bias and unpredictably skews the shape bias by a few percent in each way as can be seen in Qwen-VL, LLaVA v1.5/NeXT/RLHF, or InternVL. Similarly, the overall largest models do not have the highest shape bias. However, scale usually improves accuracy.
Does RLHF align shape bias?
RLHF-tuned VLMs are still rare at this point and we only have three samples. On both LLaVA-RLHF [69] models we see no changes in comparison to the default LLaVA models. GPT-4V [12] (though it is unclear if vision was also RLHF trained) shows one of the lowest shape biases in our study, but we do not know how the base model ranks. Overall it is hard to derive a conclusive answer, but it seems that at least RLHF does not necessarily guarantee an alignment of visual preferences.
4.2 Examining Decision Formation
Figure 3: VLMs provide high-confidence answers in our test that prefer one specific cue and ignore the other. Confidence only slightly differs by cue, but answers with the highest confidence exclusively belong to shape decisions. (a) Confidence distribution of shape and texture tokens for all samples, (b) Confidence distribution in correct predictions (i.e., the predicted label matches either shape or texture), (c) Evolution of shape bias if filtering for confidence. Measured on LLaVA-NeXT 7B for the VQA task.
To better understand how decisions are formed we utilize a confidence measurement as described in Sec. 3.2. Exemplarily, we choose LLaVA-NeXT 7B [62] to study the sampling of responses in VQA. The choice of this model is motivated by the fact that we can obtain access to the logits, in addition to the model’s ability to consistently follow the instructions as opposed to only responding with the option letter.
@@ 4.2 Examining Decision Formation
In Fig. 2(a), we visualize the confidence of the token corresponding to the shape or texture answer option. We were expecting some normal distribution that is slightly offset in favor of shape bias. However, to our surprise, we find that confidence in both options is almost binary. Analogously, when we only focus on correct answers (Fig. 2(b)), we observe that the model is highly confident in its responses. As the model assigns such a high confidence to the cue it answers with, this suggests that information from the other cue, in fact, gets lost in processing. This is also backed by the fact that the second (significantly less confident) prediction token often ($70.7\%$) does not belong to the conflicting cue. Only $17.7\%$ of the top-2 pairs contain both, shape and texture.
On average, we also observe that the model is slightly less confident in texture decisions ($\tilde{p}=0.87$ vs. $0.92$). This becomes even more evident when we only analyze responses above a confidence threshold (Fig. 2(c)). Up until $p=0.5$, this does not significantly influence any of the measurements. Beyond this point, the ratio of shape decisions increases without a decrease in texture decisions resulting in a higher shape bias. Beyond $p=0.9$ the ratios diverge, and the model reasons exclusively on shape cues.
It is not clear if these findings generalize to other models - e.g., GPT-4 was shown to be significantly less confident compared to its base model [70]. However, LLaVA is currently the building block for many successors [62, 69] and as such will strongly induce its bias into those. The next section also shows that many VLMs actually reason very similarly in our tasks hinting that our results may generalize after all.
4.3 The Role of the Vision Encoder
The texture/shape bias is primarily a vision bias - we ablate the vision encoder to understand how it drives the model decision. For this section, we additionally dissect inference from encoder pretraining.
Do VLMs just inherit the shape bias from their encoders?
Most VLMs combine a (frozen) CLIP vision tower with an LLM via some projector [23, 21, 13]. This raises the question of whether these VLMs simply learn to perform zero-shot classification using their encoders whenever the prompt requires some form of classification - thus, inheriting the shape bias from the encoder. To gain insights into this we measure the respective difference in shape bias (and accuracy).
We see that shape bias is indeed (strongly) influenced by the encoder, but the VLM can deviate in either direction ($-9.5\%$ to $+7.4\%$ in our examples in the appendix). Further, all models decrease in accuracy compared to their encoders (all of which were frozen during expansion into VLMs). This is not surprising, as VLMs are less specialized, but it shows that VLMs do not simply pass through their encoder’s classifications and have some form of bias control.
As an additional signal, we measure the error consistency [71] for insights into the alignment: a pair of models with high error consistency will share some traits in their internal decision-making process. We observe that VLMs show a high but not perfect error consistency to their respective encoders (up to $79.8\%$). Generally, they are fairly similar to all zero-shot classifiers that we benchmarked (except CNN-based models; see below). This trend even prevails in models with different (MoE)-LLMs but the same encoder. In line with our previous results, this further emphasizes that shape bias in VLMs is influenced by the encoder, but their multimodal fusion with LLMs allows them to deviate beyond that.
Does multimodal training guarantee shape-biased encoders?
Next, we aim to understand a critical aspect of encoder training: the combination of vision and language. It was already demonstrated that ViT-CLIP models show an increased shape bias in comparison to vision-only models independent of their architecture, training data, or method [29]. This may suggest that joint embedding alone increases shape bias. To verify this hypothesis, we measure the shape bias of a ResNet-50-based CLIP [11] - and observe an opposite trend. With just $20.8\%$ (refer to the appendix for details), this model even slightly reduced shape bias compared to an ImageNet-trained ResNet-50 ($22.2\%$). As such, just fusing language into encoder training does not guarantee an increased shape bias and the results strongly depend on the vision architecture, as well. For the design of shape-biased VLMs it is, thus, reasonable to rely on representations of ViT-based CLIP as opposed to vision-only models.
Figure 4: VLMs make similar errors on the cue-conflict datasets and share similarities with their vision encoders. In terms of errors, VLMs are also more similar to humans than ImageNet-trained/finetuned models. We measure the pair-wise error consistency [71] between predictions. For this analysis, an error is any answer that does not belong to the shape class (analogous to [29]). Shown responses belong to LLM-based VLMs (under the VQA task), other selected models including ImageNet models, (some) VLM encoders under ImageNet-finetuning and zero-shot classification, and ten human subjects.
5 Steering Texture/Shape Bias
@@
In the previous section, we have seen that in VLMs, visual biases are not simply inherited from the vision encoder, but the fusion with an LLM (including training) plays a crucial role. We here seek to test how flexible the representation of texture/shape bias is and whether we can systematically steer it - first through visual steering or preprocessing (Sec. 5.1), and then through language or prompt engineering (Sec. 5.2). Since texture/shape bias is a visual bias, we generally expect more flexibility with vision steering. However, being able to steer a visual bias through language alone provides a powerful way of shaping visual biases in a user-specified way.
5.1 Steering in Vision
Figure 5: Image preprocessing can strongly steer texture/shape bias. Left: Shuffling image patches with decreasing patch size results in a strong texture bias, Right: Increasing Gaussian noise introduces a strong shape bias.
Earlier work demonstrated that ImageNet-models can still detect objects even if the image is split into patches and shuffled [72, 6, 9]. As patch size decreases, the operation is destroying more global shape information, yet retaining local texture information. While it is straightforward to assume that this also applies to VLMs and significantly reduces shape bias, we emphasize quantifying the shape bias precisely and understanding whether detection is still possible for these models, as they might be more intricate than specialized models.
Oppositely, to increase the shape bias we experiment with Gaussian noise processed inputs. This is loosely inspired by applying “diffusion-like noise” during training and inference which has been shown to drastically improve the shape bias of ImageNet-ResNets [31]. However, we only apply the noise during inference and use a more simplistic approach by adding $\mathcal{N}(0,\sigma^{2})$ noise to all channels, consecutively clamping values to $[0,1]$.
We show results on LLaVA-NeXT 7B in Fig. 5, also in comparison to its CLIP encoder under zero-shot classification, and finetuned on ImageNet. Adding noise results increases the VLMs shape bias up to $89.5\%$ at $\sigma^{2}=0.5$, and patch shuffling decreases shape bias (increases texture bias) to $8.4\%$ at $28\times 28$ patches. We see similar trends for the encoder, but the VLM overall achieves the highest shape bias. Inspired by these strong results, we repeat the experiments on the more shape-biased and larger InternVL-Chat 1.1 [2]. In this model, we can further extend the range $91.7\%$ shape bias ($\sigma^{2}=0.3$), and down to $6.1\%$ ($28\times 28$ patches). While extreme steering of shape bias comes at a cost in accuracy, all results are still well beyond random chance (min: 48.6% accuracy).
In conclusion, the shape bias can be strongly steered in the vision representation (including processing by the LLM) to almost both extremes—more shape-biased and more texture-biased.
5.2 Steering in Language: Prompt Engineering
Our previous results suggest that VLMs learn a connected multimodal understanding of shape and texture. This opens the question of whether visual biases can be influenced through text processing in these models. We test this hypothesis by recording texture/shape bias as a function steering them via text through prompt engineering.
Bias steering through hand-crafted prompting.
We start by asking VLMs to specifically identify either the “shape” or the “texture” category in a given cue-conflict image. This hand-crafted biased prompting approach does indeed steer shape bias to a certain degree: As shown in Fig. 6, prompting can steer a visual bias (without significantly affecting accuracy). Details on these biased prompts can be found in the appendix. Neutral prompts perform often similarly to shape-biased prompts, whereas texture-biased prompts deviate more significantly. This suggests that models may be more inclined to use shape by default, but also have access to a certain amount of texture information which can be accessed through biased prompting.
For two models (InternVL-Chat 1.1 [2] and LLaVA-NeXT 7B [62]) we additionally replace the terms texture/shape by strong synonyms obtained from Thesaurus.com [73, 74]. Synonyms of either term can steer shape bias as well to a certain degree. For “texture” synonyms we observe more variance, as “texture” is overloaded by different meanings (e.g., some synonyms like “feeling”, “taste”, or “touch” are unrelated to texture in vision). In contrast, “shape” is a fairly well-defined term.
While the effect of steering by language is systematically visible, language steering alone does not fundamentally change the reliance on the underlying cue. This effect does not appear to be a limitation of LLM capacity: We performed an additional study on InternVL-Chat 1.2+ (34B vs. 13B) but did not obtain evidence that larger LLMs guarantee more steerability.
@@ 5.2 Steering in Language: Prompt Engineering
Figure 6: Prompts can steer the texture/shape bias to some extent. We test the same texture/shape-biased instructions on multiple models, showing that these can already shift some decisions (usually in favor of texture). For InternVL 1.1 and LLaVA-NeXT 7B we additionally test the understanding of texture/shape by using synonyms. Furthermore, we use an LLM to automatically search for specific prompts to optimize in either direction.
Interestingly, our findings of steering by language show parallels to a human psychophysical laboratory experiment. [1] conducted control experiments in which humans were either asked to identify the shape while ignoring texture, or conversely to identify the texture while ignoring the shape. This “human prompt steering” worked, but only to a certain extent: When humans were tasked to ignore the shape, the human shape bias decreased from $96\%$ (neutral instruction) only to approx. $70\%$ shape bias (texture-biased instruction). Our models behave somewhat similarly: their visual shape bias can be steered through prompting, but it appears hard for them to completely go against their default visual bias.
Bias steering through automated prompt engineering.
Above we observed that hand-crafted prompts can steer visual biases, but only to a limited extent. Does this indicate a limit on how much language/prompting can influence biases, or merely reflect that the handcrafted prompts were chosen suboptimally? To investigate this question, we tested an automatically crafted prompt. This is achieved by employing an LLM as optimizer [45] to continuously generate new prompts in natural language targeting to maximize either shape or texture bias in a feedback loop. We provide the LLM feedback about the achieved accuracy and shape bias. Additionally, we opt for greedy token sampling in the VLM to reduce noise in the feedback loop. For further details, we refer the reader to the appendix.
@@
The results are shown in Fig. 6 in gray and denoted as “search”. We observe that for both LLaVA-NeXT-7B and InternVL-Chat 1.1, automatically generated prompts exceed the manually crafted biased prompts in terms of their effectiveness to increase texture bias, and roughly match them when it comes to increasing shape bias. For InternVL-Chat 1.1 the delta between both extremes is $23.3\%$, which can only serve as a lower bound and is likely improvable by a better design of the LLM task (or using other optimizers). In line with hand-crafted prompts, overall accuracy does not change considerably, which is a key differentiator from the vision-steering we explored in the section above. We should also note that the optimization here is done with respect to the cue-conflict test set; this is simply done as a proof of concept to show that there are prompts which are able to substantially influence visual biases and not to claim a SOTA shape bias. It would be interesting to test whether the generated prompts generalize to other models or datasets.
Taken together, language-based steering clearly has a more reduced effect compared to visual steering. That said, unlike visual steering through image modifications which substantially hurts accuracy, our prompting strategy barely affects accuracy. Additionally, the method may be more intuitive for practitioners who can access or change the prompt as desired.

### v2

@@ 4.1 Key Results
The results in Fig. 2 paint a fairly uniform picture across different models and on two different tasks. Overall, the shape bias of VLMs is still significantly lower than that of humans ($96\%$), but higher than in typical image-only discriminative classifiers (e.g., $22\%$ for an ImageNet-trained ResNet-50 He et al. (2015); Geirhos et al. (2019)). Additionally, for most models, the shape bias is higher than the ca. 60% shape bias of CLIP ViT-L/14 Radford et al. (2021) - an interesting result given that this model is a common vision encoder used in many of our tested models. GPT-4V OpenAI (2023) is an unexpected outlier both in terms of accuracy and in terms of texture bias. We further discuss this particularity in Sec. 6.
The task only marginally affects the shape bias. Despite conceptually different tasks, i.e., the discriminative VQA task and the one open-ended captioning task, we do not observe fundamental shifts in the utilized information cue. We were able to report shape bias under the image captioning task for all models. However, a few models did not follow the VQA instructions and are, thus, not reported in Fig. 2. Most of these models displayed a pronounced texture bias, which might hint towards a correlation between underfitting and texture bias, but to answer this question conclusively, we would need more samples.
On average, the shape bias is slightly higher for the image captioning task than for VQA (on average 63.9% versus 61.6% for those models that could be evaluated on both tasks). However, this comes at some cost in accuracy (on average 71.0% versus 78.9%). This decrease in accuracy is due to generic captions that do not refer to any class (see Appendix B for details). For VQA the range is from $52.9$ $73.8\%$ and $54.1$ $73.2\%$ for captioning - yet, outliers with a significantly lower ($38.2\%$) shape bias in captioning exist and for the individual models, the cue bias strongly depends on the considered task (refer to the gap between circles and stars for several of the models in Fig. 2). Exceptions seem to be, for example, the Gemini Pro Vision 1.0 model, several of the LLaVA models, and the InternVL-Chat 1.2+ model for which the considered task barely influences the cue bias.
4.2 Mechanistic Analysis
We have observed that the shape bias in VLMs differs from that of CLIP (ViT-L/14), the vision encoding model used in most of the tested models. This prompts an inquiry into the VLM’s decision process. Specifically, it raises the question of whether language can affect the purely visual texture/shape bias. In this section, we want to specifically look at the vision encoder and its representation having only access to the visual input, and the LLM combining both modalities.
Table 1: Comparison between VLMs and their encoders. We show the relative difference between a VLM and its encoder in shape bias and accuracy when evaluated on cue-conflict tasks, along with error consistency Geirhos et al. (2020b). VLM performances are assessed using VQA, while the vision encoders are evaluated using zero-shot classification. Statistically significant changes in shape bias ($p<0.05$ in a two-sided t-test) are denoted by an next to the value.
VLM - Encoder [%] Error Con-
VLM Vision Encoder Accuracy Shape Bias sistency [%]
LLaVA v1.5 13B Liu et al. (2023a)  CLIP ViT-L/14@336px Radford et al. (2021) $-3.50$ $+4.2$ 73.5
LLaVA v1.5 7B Liu et al. (2023a) $-3.00$ $+1.5$ 76.8
LLaVA-NeXT 34B Liu et al. (2024) $-9.92$ $-3.9$ 67.1
LLaVA-NeXT 13B Liu et al. (2024) $-0.33$ $-2.7$ 70.9
LLaVA-NeXT 7B Liu et al. (2024) $-1.08$ $-0.2$ 67.5
MoE-LLaVA v1.5 Phi2 x4 Lin et al. (2024) $-1.42$ $-0.3$ 73.4
MoE-LLaVA v1.5 Qwen x4 Lin et al. (2024) $-24.25$ $+3.0$ 51.4
MoE-LLaVA v1.5 StableLM x4 Lin et al. (2024) $-3.67$ $-0.8$ 73.1
InstructBLIP FLAN-T5-XL Dai et al. (2023)  EVA-01-CLIP ViT-g/14@224px Sun et al. (2023b) $-6.83$ $+1.8$ 73.7
InstructBLIP Vicuna-7B Dai et al. (2023) $-14.42$ $+7.4$ 78.7
Emu2-Chat Sun et al. (2023a)  EVA-02-CLIP-E/14+@448px Sun et al. (2023b) $-11.08$ $-9.5$ 61.0
4.2.1 Vision Encoding
Most VLMs combine a frozen CLIP vision tower with an LLM via some projector Dai et al. (2023); Liu et al. (2023b); Alayrac et al. (2022). Hypothetically, the LLM could learn to perform zero-shot classification using their encoders akin to a function call whenever the prompt requires some form of classification and then simply forward the result. In such a case, the VLM would also inherit the shape bias from the encoder. To gain more insights, we ablate the encoder using zero-shot classification from the full model. We derive the encoder’s predictions by calculating the cosine similarity between the encoded class labels [^2] We explore various prompt templates (and ensembles) in Appendix J, yielding consistent results. and the input sample, selecting the label with the highest similarity to the image. Specifically, we measure the difference in accuracy and shape bias between the encoders of the VLM and the VLM itself, as well as their error consistency Geirhos et al. (2020b).
Error consistency is a metric to assess whether two observers (e.g., a human and a model) systematically make errors on the same images. If that is the case, this suggests a deeper underlying similarity compared to simply reaching similar overall accuracies, since those could be reached with very different strategies. The metric is based on Cohen’s kappa Cohen (1960) and computes how frequently errors made by two observers overlap (up to perfect overlap) while correcting for the overlap expected by chance. Cohen’s kappa is within $[-1,1]$, with a value of 00 indicating chance-level consistency, positive values indicating systematic agreement, and negative values indicating systematic disagreement. For our measurement, we treat all predictions other than the shape label that are not related to the shape cues as errors.
The vision encoder provides a flexible representation. Our comparisons in Tab. 1 show that VLMs’ decisions differ from their isolated encoders. Even on the rather simple 16-way cue-conflict problem, all VLMs decrease in accuracy compared to their encoders in zero-shot classification. This may be expected as the increase in supported tasks in VLMs comes at a cost in specialization, but is already a sign of changes. Further, we note that the error consistency to their respective encoders only matches up to $78.7\%$. This leaves at least $20\%$ room for decisions that differ from the encoder, proving that the LLM and the text prompt further influence the shape bias, despite being a vision-only bias. This deviation can only be possible if the generated vision tokens are flexible to some degree by containing information belonging to both cues. This is further confirmed by the measurement of shape bias, which shows a $-9.5\%$ to $+7.4\%$ difference in both directions.
4.2.2 LLM Processing of Vision Tokens
((a)) LLaVA-NeXT 7B
((b)) InternVL-Chat 1.1
((c)) MoE-LLaVA-Phi2
Figure 3: Confidence distribution of shape and texture tokens for all samples. All models form highly biased decisions by completely ignoring one cue. Measured on LLaVA-NeXT 7B, InternVL-Chat 1.1, and MoE-LLaVA-Phi2 for the VQA task.
In the VQA task, we force the model to predict a single class - yet, the previous section has shown that the vision tokens generated by the encoder contain information from both cues. To better understand the processing, we evaluate the VQA prediction confidences as follows.
All answer options in our VQA prompts correspond to a single character and, thus, a token. Well-behaving models, where the response consistently starts with the option letter (and nothing else) allow us to gather insights into the prediction process. For these models, the logits of each token correspond to the logits of option letters. By applying a softmax function, we can analyze the confidence in each option (we map invalid tokens to a separate null class). This allows us to better study the sampling behavior and make the following observation:
LLMs turn flexible representations into biased decisions. In Fig. 3, we visualize the confidence of the token corresponding to the shape or texture answer option. This experiment can only be performed on models where we have access to the logits, and the model consistently follows the instructions to only respond with the predicted options letter. This limits the analysis to a few models in our zoo, and we show similar results on LLaVA-NeXT 7B Liu et al. (2024), InternVL 1.1 Chen et al. (2024) and MoE-LLaVA-Phi2 Lin et al. (2024).
@@ 4.2.2 LLM Processing of Vision Tokens
To our surprise, we find that confidence in both options is almost binary. Analogously, when we only focus on correct answers, we observe that the model is highly confident in its responses. As the model places such high confidence in the selected cue, this suggests that information from the alternative cue is effectively disregarded during LLM processing. This conclusion is further supported by the observation that the second, significantly less confident prediction token does not align with the conflicting cue. For example, in LLaVA-NeXT 7B Liu et al. (2024), this occurs in 70.7% of cases. Only $17.7\%$ of the top-2 pairs contain both, shape and texture. Thus, while the encoder has its own inductive bias that directs the final decision, the actual biasing happens in the LLM, similar to linear classification heads in discriminative models Islam et al. (2021).
While it is not clear if these findings generalize to all other VLMs, the overall high error consistency between all VLMs (see Fig. 13 for a heatmap) on our task hints that our results may generalize well to other models.
5 Prompt-based steering of VLM outputs
@@
In the previous section, we have seen that in VLMs, visual biases are not simply inherited from the vision encoder, but the fusion with an LLM including the text prompt plays a crucial role. Given the somewhat flexible representation of texture/shape bias, we test in the following if we can actively talk VLMs into seeing the world differently, i.e., systematically steer the output towards either end of the bias and go beyond the inductive bias. Therefore, we explore the influence of language through prompt engineering on the visual bias (Sec. 3.1.1). Furthermore, since texture/shape bias is a vision-only bias, we contrast this language steering to visual steering through image preprocessing (see Appendix F). While we expect strong model steerability through image preprocessing (obviously at high costs in model accuracy), Fig. 1 indicates significant flexibility of (some) VLMs through language, potentially offering a powerful way of shaping visual biases in a user-specified way without the need to retrain a model.
@@
Figure 4: Language can steer the texture/shape bias to some extent. We test the same texture/shape-biased instructions on multiple models, showing that these can already shift some decisions (usually in favor of texture). The stated percentages refer to the achieved accuracy on cue-conflict. For InternVL 1.1 and LLaVA-NeXT 7B we additionally test the understanding of texture/shape by using synonyms. Furthermore, we use an LLM to automatically search for specific prompts to optimize in either direction.
5.1 Steering towards Shape- or Texture-biased Outputs
Our previous results suggest that VLMs learn a connected multi-modal understanding of shape and texture. This opens the question of whether visual biases in outputs can be influenced through text processing in these models. We test this hypothesis by recording texture/shape bias as a function, steering it via text through prompt engineering.
Bias steering through hand-crafted prompting. We start by asking VLMs to specifically identify either the "shape" or the "texture" category in a given cue-conflict image. As shown in Fig. 4, prompting can steer a visual bias (without significantly affecting accuracy). Details on these biased prompts can be found in Sec. C.3. Neutral prompts often perform similarly to shape-biased prompts, whereas texture-biased prompts deviate more significantly. This suggests that models may be more inclined to use shape by default, but also have access to a certain amount of texture information which can be accessed through biased prompting.
To better test the representation of "shape" and "texture", we additionally replace the terms texture/shape with strong synonyms obtained from Thesaurus.com Dictionary.com (2024a; b). Then we measure shape bias on (InternVL-Chat 1.1 Chen et al. (2024) and LLaVA-NeXT 7B Liu et al. (2024)). Synonyms of either term can steer shape bias as well to a certain degree. For "texture" synonyms, we observe more variance, as "texture" is overloaded by different meanings (e.g., some synonyms like "feeling", "taste", or "touch" are unrelated to texture in vision). In contrast, "shape" is a fairly well-defined term. This demonstrates that the steering is not coincidental but leverages a learned representation.
While the effect of steering by language is systematically visible, language steering alone does not fundamentally change the reliance on the underlying cue. This effect does not appear to be a limitation of LLM capacity - the evaluation of InternVL-Chat 1.2+ (34B vs. 13B) does not provide evidence that larger LLMs offer more steerability.
Bias steering through automated prompt engineering. Did our results indicate a limit on how much language/prompting can influence biases, or merely reflect that the handcrafted prompts were chosen suboptimally? To address this question, we test automatically crafted prompts (see Sec. 3).
@@ 5.1 Steering towards Shape- or Texture-biased Outputs
The results are shown in Fig. 4 in gray and denoted as "search". We observe that for both LLaVA-NeXT-7B and InternVL-Chat 1.1, automatically generated prompts exceed the manually crafted biased prompts in terms of their effectiveness to increase texture bias, and roughly match them when it comes to increasing shape bias. For InternVL-Chat 1.1 the delta between both extremes is $23.3\%$, which can only serve as a lower bound and is likely improvable by a better design of the LLM task (or using other optimizers). In line with hand-crafted prompts, overall accuracy does not change considerably or sometimes even improves. We should also note that the optimization is done for the cue-conflict test set; this is simply done as a proof of concept to show that there are prompts that can influence visual biases substantially, and not to claim a SOTA shape bias.
5.2 Steering towards Low- or High-Frequency-biased Outputs
Table 2: Prompt steering on frequency-cue-conflict. Statistically significant changes are marked by * (two-sided t-test with $p<0.05$). We compare the "neutral" prompt with found prompts to maximize ("search (max)") or minimize the respective bias ("search (min)").
Model Prompt Accuracy [%] LF Bias [%]
InternVL-Chat 1.1 neutral 92.92 34.5
search (max) 90.33 38.6*
search (min) 91.33 32.9
LLaVA-NeXT 7B neutral 82.83 52.4
search (max) 84.25 54.5
search (min) 82.67 48.7*
So far, we have analyzed the texture/shape bias. In this section, we show that steering is possible for other biases, too. To this end, we explore a bias originating in the spectral domain, specifically focusing on low versus high-frequency cue conflicts (Oliva et al., 2006). This bias has been shown to affect a classification model’s robustness, for example in Wang et al. (2020); Lukasik et al. (2023). A related observation has been made in Subramanian et al. (2024), showing that the critical frequency band of object recognition separates human perception from model vision.
To study the steerability of the low-frequency bias (LF bias), we propose a new dataset of stimuli following the texture/shape cue-conflict benchmark methodology Geirhos et al. (2019): We create 1,200 samples belonging to 16 ImageNet-super-categories by blending the spectral components of two differently labeled images: 30% of the low-frequency components from one image and 70% of the high-frequency components from the other. Please find the details Appendix G.
We test the same neutral prompt as for previous experiments and use our automated prompt search to either maximize or minimize the bias. The results in Tab. 2 show that again we can steer the bias by language. The range of course depends on the vision representation and language training and is not as pronounced as for texture/shape bias. Still, we find that our prompts can result in statistically significant changes in bias. On LLaVA-NeXT 7B, the prompts even improved accuracy alongside the bias.
Independent of the prompting, it is interesting that the smaller LLaVA-NeXT 7B model almost perfectly balances the conflicting cues, whereas the larger InternVL 1.1 model is significantly biased toward HF. We hope that this dataset can pave a new avenue for future research on frequency bias.